OpenStudy (anonymous):

In a mixture of concrete there are two lb of cement mix for each pound of gravel. If the mixture contains a total of 54lbs of these two ingrediants, how many pounds of gravel are there?

OpenStudy (perl):

Solve the proportion : 2 lb cement / 1 lb gravel = (54 - x) / x

OpenStudy (anonymous):

I don't know how to solve these well can you show me how

OpenStudy (perl):

solve 2/1 = (54 - x ) / x you can cross multiply

OpenStudy (anonymous):

how do you cross multiply?

OpenStudy (perl):

2 * x = 1 * ( 54 - x )

OpenStudy (perl):

2x = 54 - x 3x = 54 x = 18

OpenStudy (perl):

so in a 54 pound mixture of concrete there are 18 pounds of gravel and 36 pounds of cement

OpenStudy (anonymous):

thank u

OpenStudy (perl):

sorry i couldnt draw it out, the draw button is broken

OpenStudy (anonymous):

its ok thnks